• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name V. Sulkosky

    Log entry time 08:01:05 on January30,2002

    Entry number 81298

    keyword=Owl Shift Summary

    00:00 Continue with 4B kinematics

    23:56-00:25 run # 1869, 4B(e) LH2-Rad Mon, Veto off, 20 uA
    00:25 Target IOC reboot, frozen screen
    00:30-01:20 run # 1870, 4B(e) LH2-Rad Mon, Veto off, 20 uA
    01:20 During run # 1870 CODA lost contact with ROC 14. CODA was killed, and the crates were reset.
    01:23-02:10 run # 1871 4B(e) LH2-Rad Mon, Veto off, 20 uA
    02:10 During run # 1871 CODA lost contact with TS2. CODA was killed, and the crates reset.
    02:15-02:24 run # 1872 4B(e) same conditions as previous runs (1869-1871)

    02:25 Magnet turned off.
    02:30 Problems with TS2 again. Restarted CODA. Run # 1873 is junk.
    02:33-03:32 Runs # 1874-1875 4B(f) LH2-Rad Moff, Veto off, 20 uA

    03:33 Magnet turned on.
    03:34-04:01 Run # 1876 4B(e) LH2-Rad Mon, Veto off, 20 uA.
    04:02 Turn Veto on.
    04:04 Asked MCC to reduce beam to 5.0 uA to reduce Veto v12 to 2 MHz.
    04:13-05:41 Runs # 1877-1879 4B(e*) LH2-Rad Mon, Veto on, 5 uA

    05:42 Asked MCC to mask the target to move to the Optics/carbon target. Turned magnet off. Veto turned off.
    05:51 Asked MCC to deliver 50 uA unrastered beam to finish 4B(g) kinematics
    06:00-06:09 Run # 1880 4B(g) Optics target Moff, Veto off, ~45 uA

    06:10 Called MCC for Controlled Access to move the calorimeter for Kinematics 4D.
    06:30 Change Left spectrometer momentum to 2.970 GeV/c.
    06:30-07:00 Access to change the calorimeter position.
    07:10 Hall A in Beam permit
    07:12 Left spectrometer angle moved at 28.16 degrees.

    07:14 Asked MCC to send 50 uA unrastered beam on Carbon target.
    07:45 I talked to MCC, and they are working on a Klystron that has been giving them problems. They say they should be up shortly.
    07:59 MCC called. They said it will be another 45 minutes to get beam back.